# Bolt3D: Generating 3D Scenes in Seconds

### v1

@@ Abstract
We present a latent diffusion model for fast feed-forward 3D scene generation. Given one or more images, our model Bolt3D directly samples a 3D scene representation in less than seven seconds on a single GPU. We achieve this by leveraging powerful and scalable existing 2D diffusion network architectures to produce consistent high-fidelity 3D scene representations. To train this model, we create a large-scale multiview-consistent dataset of 3D geometry and appearance by applying state-of-the-art dense 3D reconstruction techniques to existing multiview image datasets. Compared to prior multiview generative models that require per-scene optimization for 3D reconstruction, Bolt3D reduces the inference cost by a factor of up to 300$\times$. Project website: szymanowiczs.github.io/bolt3d.
Figure 1: Given an arbitrary number of input images, Bolt3D directly outputs a 3D representation which can be rendered at interactive frame-rates. Operating in a feed-forward manner, generation takes mere seconds. Bolt3D features a latent diffusion model with a scalable 2D architecture, trained on large-scale appearance and geometry data, enabling generation of full 360 scenes from one or multiple input images. We encourage the reader to explore these scenes in the interactive viewer available on the project website.
@@ 1 Introduction
We demonstrate that Bolt3D outperforms prior single- and few-view feed-forward 3D regression methods and synthesizes detailed 3D scene content, even in ambiguous regions that are not observed in any input image. Furthermore, we show that Bolt3D reduces inference cost up to 300×300$\times$300 × compared to multi-view image generation methods, which typically require per-scene optimization to reconstruct 3D models.
@@ 2 Related work
Reconstructing a 3D scene from an input image (or from multiple images) is sometimes formulated as a feed-forward regression problem. Many different 3D representations have been developed towards this goal, including multi-plane images [54, 25], voxels [51, 53], meshes [15, 13], and radiance fields [66, 20]. Most relevant to us are recent approaches that output 3D Gaussians [23], a representation favored for its real-time rendering speed. Splatter Image [49] and pixelSplat [4] were the first to associate Gaussians with the pixels of one or two input images, respectively. Follow-up works iterated on this approach with improved architectures [73], depth conditioning [50, 63], explicit feature matching [5], and removing camera pose requirements [65]. Although these regression-based methods are capable of accurately reconstructing observed regions, they tend to produce blurry results in unseen regions. In contrast, Bolt3D is a generative approach and is thereby capable of generating unobserved regions of the scene.
@@
A common paradigm for addressing ambiguity in few-view 3D reconstruction is to generate multiple views of the scene, from which 3D can be recovered. Zero-1-to-3 [31] and 3DiM [59] did this for individual objects, and follow-up work improved the conditioning mechanism in a variety of ways [3, 61, 16, 19, 78, 44]. Joint sampling improves sample consistency, and can be achieved by single-image [26], multi-view generators [12, 56, 21, 32], or video models [34, 55]. Alternatively, other methods have explored using estimated depth to improve geometric consistency [33, 21, 70, 68] The outputs of the generation process are often fed into a reconstruction pipeline [36, 23, 58]. This two-stage approach produces high-quality results, but requires minutes or hours of optimization time for real-world scenes. Our approach leverages a diffusion model to generate multiple views of the scene, but in our approach the 3D scene is a direct output from our model, so no optimization stage is required and the cost of inference is therefore decreased by multiple orders of magnitude.
@@
Directly generating 3D representations circumvents the limitations of the previous works, by combining the ability of generative models to handle ambiguity while avoiding a costly distillation/optimization stage. Early diffusion-based models denoise a voxel grid that parameterizes a radiance field [37] require 3D supervision. The difficulty in obtaining such 3D data motivated the formulation of a denoising objective on 2D images with a rendering bottleneck [48, 52], similar to early GAN [14]-based works that directly output voxel grids [17], radiance fields [45, 69] or tri-planes [2] using differentiable rendering to directly train on 2D images. Follow-up works improve the architecture [64] and the 3D representation [57, 28, 75], but such approaches have shown limited success beyond individual objects and small baseline camera motion. Some more recent works [27, 67, 6, 75] exploit temporal locality to generate videos given camera trajectories. A second body of work aims to improve the efficiency of the generation process by using fine-tuning pretrained VAEs to reduce the computational burden of the generation process, showing remarkable success on bounded objects [28, 35, 41, 62, 74, 72]. We similarly train a VAE on a joint image-geometry latent space, but our focus is on complex, unbounded scenes. Most similar to our work is latentSplat [60] which learns a latent geometry representation similar to ours. However, latentSplat is limited to single object categories, is only applicable to 2-view reconstruction and features a VAE-GAN framework. Our model works on any object or scene, can take any number of images as input (as few as 1) and builds on a powerful latent diffusion model. We also demonstrate that our method performs better.
@@ 3 Preliminaries
Other parametrizations of the denoiser model have been proposed, such as predicting the noise added to the clean data or a combination of the clean data and the noise like the ${\mathbf{v}}$-prediction [43]. A latent diffusion model first compresses the data into a lower-dimensional latent space and then builds a diffusion model in the latent space. A two-stage training procedure is typically applied: the compression is first learned by a VAE framework and fixed, followed by learning the diffusion model.
@@ 4.2 Geometric multi-view latent diffusion model
We finetuned our model from a pretrained multi-view image diffusion model, which itself was finetuned from text-to-image latent diffusion model, to maximally maintain the generalization ability of the model while being trained on limited amount of multi-view data with 3D positions. The images and geometry (i.e., pointmaps and cameras) are encoded and decoded by two separate VAEs that 8×8$\times$8 × downsample the input signals spatially (i.e. from $512\times 512$ in the pixel space to $64\times 64$ in the latent space). The image VAE is pre-trained and frozen while the geometry VAE is trained from scratch. The camera pose is parametrized as a $6$-dim raymap that encodes the ray origin and direction at each spatial location. At the input of the latent diffusion, the image latent, geometry latent and a raymap of the same size of the latents are channel-wise concatenated. We use a ${\mathbf{v}}$-parametrization and a ${\mathbf{v}}$-prediction loss for the diffusion model [43]. The model is pre-trained on $8$ and fine-tuned on 16 views in total, and $1$-$3$ views are randomly sampled as input views. During sampling we generate $16$ views in total.
@@ 4.3 Gaussian Head
Given cameras and generated images and pointmaps, we train a multi-view feedforward Gaussian head model to output the refined color, opacities and covariance matrices of 3D Gaussians stored in Splatter Images. We first calibrate the generated pointmaps to be pixel aligned under the corresponding camera coordinate system. That is, we transform the 3D points into the camera coordinate, keep $z$ and set $x,y$ based on the camera ray and $z$. While the VAE decodes color and pointmaps independently for each view, we found that implementing a multi-view Gaussian head model is crucial. A U-ViT architecture is applied with 4×4$\times$4 × patchification before the transformer blocks, see Supplementary for more details. The Gaussian head takes 8 views as input, and is trained using photometric losses (i.e. an L2 loss and a perceptual loss [76]) on rendered images from $4$ subsampled input views and $8$ novel views.
@@ 4.4 Training
We adopt a 3-stage training process:
@@
    Geometry VAE. We first train the model at 256$\times$256 resolution for 3 million iterations, then finetune at 512$\times$512 for another 250k iterations.
@@ 5 Experiments
Generative methods are best evaluated qualitatively, so we begin our experiments with presenting the results of our method in Fig. 3. In addition, we encourage the reader to view the videos and interactive visualizations in the supplementary material. We present high-quality reconstructions on a wide range of inputs.
@@
We quantify 3D reconstruction quality with standard metrics for novel-view synthesis: PSNR, SSIM, and LPIPS [76] to measure pixel-wise, patch-wise and perceptual similarity respectively.
@@
We use 1-view RE10K for comparison against Flash3D by using the split from [61, 12] and using the first frame in each video as the conditioning frame, with the target frames as specified in the split. This split contains larger camera motion (90 frames) than that used in the original Flash3D paper, thus evaluating the ability to reconstruct scenes beyond small ($\pm$ 30 frames) camera motion. Additionally, we include an evaluation on CO3D, also adapted from [61, 12] in the same manner as RE10K to evaluate both methods under stronger self-occlusion, as is typical for single-view reconstruction. For comparison against DepthSplat [63], we use DL3DV, which features large camera motion, and we evaluate both methods in the 2-view and 4-view settings, using source and target views from [63]. We run evaluations on the scenes using the overlap of their and our testing sets.
@@ Results.
We find that Bolt3D outperforms both methods, as shown in Tab. 1. The strong performance of our method compared to feed-forward reconstruction suggests that modeling ambiguity is important, as we show in Fig. 4. This observation is consistent with the fact that the biggest gain in performance is observed in the 1-view setting, where ambiguity is the largest.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
1-view Flash3D 17.40 0.699 0.419
RE10K Ours 21.03 0.805 0.257
1-view Flash3D 14.43 0.552 0.608
CO3D Ours 16.78 0.562 0.505
2-view DepthSplat 16.25 0.515 0.465
DL3DV Ours 17.75 0.551 0.392
4-view DepthSplat 19.48 0.638 0.327
DL3DV Ours 20.64 0.653 0.310
Table 1: Quantitative comparison to regression-based methods. Our generative approach improves performance across multiple datasets and numbers of input views. The biggest gain is seen in the 1-view setting where ambiguity is the greatest.
@@
When comparing to LatentSplat, we evaluate at 256$\times$256 resolution on the datasets they proposed: RealEstate10k and the hydrants and teddybear categories of CO3D, with their extrapolation splits. When comparing to Wonderland, we follow their quantitative evaluation protocol on RealEstate10K, i.e. randomly sampling 1000 testing scenes, using the first frame of the video as the source frame and measuring the rendering quality of 14 views following the source view given that one input view.
@@
We observe that our method outperforms LatentSplat qualitatively (Fig. 5) and quantitatively (Tab. 2). Fig. 5 illustrates that our diffusion model generates higher-quality details than LatentSplat’s VAE-GAN. We also demonstrate that our method outperforms Wonderland in Tab. 2. We hypothesize this is due to explicit modeling of geometry in the autoencoder and in the diffusion model. It is also worth noting that Wonderland uses a video model and takes 5 minutes to generate a scene, while ours generates 16 splatter images in around 6 seconds.
Figure 5: Our latent diffusion framework generates more realistic details than the VAE-GAN model proposed by LatentSplat.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
1-view Wonderland* 17.15 0.550 0.292
Re10k Ours 21.54 0.747 0.234
2-view LatentSplat 22.62 0.777 0.196
RE10K Ours 23.13 0.806 0.166
2-view LatentSplat 15.78 0.306 0.426
CO3D hyd. Ours 17.38 0.437 0.390
2-view LatentSplat 17.71 0.533 0.434
CO3D ted. Ours 18.94 0.605 0.393
Table 2: Quantitative comparison to 3D generative methods. Our diffusion-based framework with explicit 3D supervision achieves the best performance. * denotes results reported in [27]
@@
Next, we explore the speed-quality trade-off versus a state-of-the-art optimization-based method, CAT3D [12]. We also include qualitative comparisons to RealmDreamer [46].
@@
In Tab. 3 we observe that our method shows strong performance, and it does so while requiring 300×300$\times$300 × less compute for inference. Qualitatively, our method produces very high quality reconstructions for a range of scenes, and in Fig. 6 we illustrate that our method generates better results than RealmDreamer, and sometimes sharper results than CAT3D, especially on the 3DGS variant and in backgrounds or fine details. This is due to the optimization process in CAT3D regressing to the mean in the case of inconsistent generations, especially in the less robust 3DGS optimization process. While our method does not outperform CAT3D, it is still capable of generating high-quality 3D scenes from a diverse range of inputs (as seen in the supplementary material), and we argue that a 300×300$\times$300 × reduction in inference cost well justifies a small drop in quality.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ gpu-min $\downarrow$
RE10 CAT3D 29.56 0.937 0.134 77.28
Ours 27.00 0.905 0.154 0.25
LLFF CAT3D 22.06 0.745 0.194 80.00
Ours 18.75 0.562 0.341 0.25
DTU CAT3D 19.97 0.809 0.202 72.00
Ours 18.59 0.738 0.312 0.25
CO3D CAT3D 20.85 0.673 0.329 73.60
Ours 19.41 0.628 0.416 0.25
Mip. CAT3D 16.62 0.377 0.515 73.60
Ours 15.67 0.309 0.540 0.25
Table 3: Comparison to optimization-based methods. Our method performs competitively with state-of-the-art optimization-based methods, while reducing the inference cost 300×300$\times$300 ×.
Figure 6: Qualitative comparison to optimization-based methods. The reconstruction quality of our method compares favorably with that of optimization-based methods, and sometimes even exhibits sharper details, while using 300$\times$ less compute.

### v2

@@ Abstract
[…]
Figure 1: Given an arbitrary number of input images, Bolt3D directly outputs a 3D representation which can be rendered at interactive frame-rates. Operating in a feed-forward manner, generation takes mere seconds. Bolt3D features a latent diffusion model with a scalable 2D architecture, trained on large-scale appearance and geometry data, enabling generation of full 360 scenes from one or multiple input images. We invite the reader to explore these scenes in the interactive viewer available on the project website.
@@ 1 Introduction
We demonstrate that Bolt3D outperforms prior single- and few-view feed-forward 3D regression methods and synthesizes detailed 3D scene content, even in ambiguous regions that are not observed in any input image. Furthermore, we show that Bolt3D reduces inference cost up to $300\times$ compared to multi-view image generation methods, which typically require per-scene optimization to reconstruct 3D models.
@@ 2 Related work
Reconstructing a 3D scene from an input image (or from multiple images) is sometimes formulated as a feed-forward regression problem [74, 34, 71, 73, 21, 19, 49, 9, 91, 28, 17, 83, 56, 79]. Most relevant to us are recent approaches that output 3D Gaussians [32], a representation favored for its real-time rendering speed. Splatter Image [68] and pixelSplat [8] were the first to associate Gaussians with the pixels of one or two input images, respectively. Follow-up works iterated on this approach with improved architectures [100, 90, 70], depth conditioning [69, 87], explicit feature matching [10], and removing camera pose requirements [89, 103]. Although these regression-based methods are capable of accurately reconstructing observed regions, they tend to produce blurry results in unseen regions. In contrast, Bolt3D is a generative approach and is thereby capable of generating unobserved regions of the scene.
@@
A common paradigm for addressing ambiguity in few-view 3D reconstruction is to generate multiple views of the scene, from which 3D can be recovered. Zero-1-to-3 [40] and 3DiM [81] did this for individual objects, and follow-up work improved the conditioning and sampling mechanism in a variety of ways [7, 84, 22, 27, 106, 58, 94]. Joint sampling improves sample consistency, and can be achieved by multi-view generators [18, 27, 76, 29, 41, 63, 104], or video models [11, 23, 43, 75, 39, 66]. Alternatively, other methods have explored using estimated depth to improve geometric consistency [42, 29, 96, 92, 93, 47] The outputs of the generation process are 2D images, and reconstructing a renderable 3D asset requires an optimization-based reconstruction pipeline [45, 32]. This two-stage approach produces high-quality results, but requires minutes or hours of optimization time for real-world scenes. Our approach leverages a diffusion model to generate multiple views of the scene, but in our approach the 3D scene is a direct output from our model, so no optimization stage is required and the cost of inference is therefore decreased by multiple orders of magnitude.
@@
Directly generating 3D representations circumvents the limitations of the previous works, by combining the ability of generative models to handle ambiguity while avoiding a costly distillation/optimization stage. Early diffusion-based models denoise a voxel grid that parameterizes a radiance field [46] and require 3D supervision. The difficulty in obtaining such 3D data motivated the formulation of a denoising objective on 2D images with a rendering bottleneck [1, 2, 67, 72], similar to early GAN [20]-based works that directly output voxel grids [24], radiance fields [59, 95, 48] or tri-planes [5, 6] using differentiable rendering to directly train on 2D images. Follow-up works improve the architecture [88] and the 3D representation [77, 36], but such approaches have shown limited success beyond individual objects and small baseline camera motion. Some more recent works [35, 61] exploit temporal locality to generate videos given camera trajectories. A second body of work aims to directly learn a ‘natively 3D’ latent space, showing remarkable success on bounded objects [36, 44, 54, 86, 101, 99], but lagging behind in quality when applied to 3D scenes with backgrounds [62, 53, 60, 98] due to limitations of the autoencoder. We opt for a view-centric representation and analyze the architecture and training recipe of the autoencoder to achieve very high visual fidelity. Most similar to our work is latentSplat [82] which learns a latent geometry representation similar to ours. However, latentSplat is limited to single object categories, is only applicable to 2-view reconstruction and features a VAE-GAN framework. Our model works on any object or scene, can take any number of images as input and builds on a powerful latent diffusion model. We also demonstrate that our method performs better.
@@ 3 Preliminaries
Other parametrizations of the denoiser model have been proposed, such as predicting the noise added to the clean data or a combination of the clean data and the noise like the $\mathbf{v}$}-prediction [57]. A latent diffusion model first compresses the data into a lower-dimensional latent space and then builds a diffusion model in the latent space. A two-stage training procedure is typically applied: the compression is first learned by a VAE and fixed, followed by learning the diffusion model.
@@ 4.2 Geometric multi-view latent diffusion model
We finetuned our model from a pretrained multi-view image diffusion model, which itself was finetuned from text-to-image latent diffusion model, to maximally maintain the generalization ability of the model while being trained on limited amount of multi-view data with 3D positions. The images and geometry (i.e., pointmaps and cameras) are encoded and decoded by two separate VAEs that $8\times$ downsample the input signals spatially (i.e. from $512\times 512$ in the pixel space to $64\times 64$ in the latent space). The image VAE is pre-trained and frozen while the geometry VAE is trained from scratch. The camera pose is parametrized as a $6$-dim raymap that encodes the ray origin and direction at each spatial location. At the input of the latent diffusion, the image latent, geometry latent and a raymap of the same size of the latents are channel-wise concatenated. We use a $\mathbf{v}$}-parametrization and a $\mathbf{v}$}-prediction loss for the diffusion model [57]. The model is pre-trained on $8$ and fine-tuned on 16 views in total, and 11-33 views are randomly sampled as input views. During sampling we generate $16$ views in total.
@@ 4.3 Gaussian Head
Given cameras and generated images and pointmaps, we train a multi-view feedforward Gaussian head model to output the refined color, opacities and covariance matrices of 3D Gaussians stored in Splatter Images. We first calibrate the generated pointmaps to be pixel aligned under the corresponding camera coordinate system. That is, we transform the 3D points into the camera coordinate, keep $z$ and set $x,y$ based on the camera ray and $z$. While the VAE decodes color and pointmaps independently for each view, we found that implementing a multi-view Gaussian head model is crucial. A U-ViT [3] architecture is applied with $4\times$ patchification before the transformer blocks, see Supplementary for more details. The Gaussian head takes 8 views as input, and is trained using photometric losses (i.e. an L2 loss and a perceptual loss [102]) on rendered images from $4$ subsampled input views and $8$ novel views.
@@ 4.4 Training
We train our model in 3 stages:
@@
    Geometry VAE. We first train the model at $\times256$ resolution for 3 million iterations, then finetune at $\times512$ for another 250k iterations.
@@ 5 Experiments
We begin our experiments with presenting the results of our method in Fig. 3, and we encourage the reader to view the videos and interactive visualizations on the project website which include high-quality reconstructions on a wide range of inputs.
@@
We quantify 3D reconstruction quality with standard metrics for novel-view synthesis: PSNR, SSIM, LPIPS [102] and FID [25] to measure pixel-wise, patch-wise, perceptual and distribution similarity, respectively.
@@
We use 1-view RE10K for comparison against Flash3D by using the split from [84, 18] and using the first frame in each video as the conditioning frame, with the target frames as specified in the split. This split contains larger camera motion (90 frames) than that used in the original Flash3D paper, thus evaluating the reconstruction ability beyond small ($\pm$ 30 frames) camera motion. Additionally, we include an evaluation on CO3D, also adapted from [84, 18] to evaluate both methods under stronger self-occlusion, as is typical for single-view reconstruction. For comparison against DepthSplat [87], we use DL3DV, which features large camera motion, and we evaluate both methods in the 2-view and 4-view settings, using source and target views from [87]. We run evaluations on the scenes using the overlap of their and our testing sets.
@@ Results.
We find that Bolt3D outperforms both methods, as shown in Tab. 1. The strong performance of our method compared to feed-forward reconstruction highlights the importance of modeling ambiguity, as we show in Fig. 4. This observation is consistent with the fact that the biggest gain in performance is observed in the 1-view setting, where ambiguity is the largest.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
1-view Flash3D 17.40 0.699 0.419 96.9
RE10K Ours 21.03 0.805 0.257 55.5
1-view Flash3D 14.43 0.552 0.608 174.8
CO3D Ours 16.78 0.562 0.505 97.5
2-view DepthSplat 16.25 0.515 0.465 95.9
DL3DV Ours 17.75 0.551 0.392 64.5
4-view DepthSplat 19.48 0.638 0.327 58.8
DL3DV Ours 20.64 0.653 0.310 48.2
Table 1: Comparison to regression-based methods. Our generative approach improves performance across multiple datasets and numbers of input views. The biggest gain is seen in the 1-view setting where ambiguity is the greatest.
@@
When comparing to LatentSplat, we evaluate at $\times256$ resolution on the datasets they proposed: RealEstate10k and the hydrants and teddybear categories of CO3D, with their extrapolation splits. When comparing to Wonderland, we follow their quantitative evaluation protocol on RealEstate10K, i.e. randomly sampling 1000 testing scenes, using the first frame of the video as the source frame and measuring the rendering quality of 14 views following the source view given that one input view.
@@
We observe that our method outperforms LatentSplat qualitatively (Fig. 5) and quantitatively (Tab. 2) on image-level metrics. We suspect that latentsplat’s adversarial loss gives it an advantage in the distribution-level FID metric. Fig. 5 illustrates that our diffusion model generates higher-quality details than LatentSplat’s VAE-GAN. We also demonstrate that our method outperforms Wonderland in Tab. 2. We hypothesize this is due to explicit modeling of geometry in the autoencoder and in the diffusion model. It is also worth noting that Wonderland uses a video model and takes 5 minutes to generate a scene, while ours generates 16 splatter images in around 6 seconds.
Figure 5: Our latent diffusion framework generates more realistic details the VAE-GAN framework proposed by LatentSplat.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
1-view Wonderland* 17.15 0.550 0.292 -
Re10k Ours 21.54 0.747 0.234 11.30
2-view LatentSplat 22.62 0.777 0.196 2.79
RE10K Ours 23.13 0.806 0.166 3.66
2-view LatentSplat 17.71 0.533 0.434 71.12
CO3D ted. Ours 18.94 0.605 0.393 76.24
2-view LatentSplat 15.78 0.306 0.426 48.03
CO3D hyd. Ours 17.38 0.437 0.390 76.12
Table 2: Comparison to 3D generative methods. Our diffusion-based framework with explicit 3D supervision achieves the best image-level metrics. * denotes results reported in [35]
@@
We explore the speed-quality trade-off versus a state-of-the-art optimization-based method, CAT3D [18]. We also include qualitative comparisons to RealmDreamer [64].
@@
In Tab. 3 we observe that our method shows strong performance, and it does so while requiring $300\times$ less compute for inference. Qualitatively, our method produces very high quality reconstructions for a range of scenes, and in Fig. 6 we illustrate that our method generates better results than RealmDreamer, and sometimes sharper results than CAT3D, especially on the 3DGS variant and in backgrounds or fine details. This is due to the optimization process in CAT3D regressing to the mean in the case of inconsistent generations, especially in the less robust 3DGS optimization process. While our method does not outperform CAT3D, it is still capable of generating high-quality 3D scenes from a diverse range of inputs (as seen in the supplementary material), and we argue that a $300\times$ reduction in inference cost well justifies a small drop in quality.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ gpu-min $\downarrow$
RE10 CAT3D 29.56 0.937 0.134 13.75 77.28
Ours 27.00 0.905 0.154 27.40 0.25
LLFF CAT3D 22.06 0.745 0.194 37.54 80.00
Ours 18.75 0.562 0.341 96.61 0.25
DTU CAT3D 19.97 0.809 0.202 41.76 72.00
Ours 18.59 0.738 0.312 67.49 0.25
CO3D CAT3D 20.85 0.673 0.329 44.33 73.60
Ours 19.41 0.628 0.416 79.60 0.25
Mip. CAT3D 16.62 0.377 0.515 91.82 73.60
Ours 15.67 0.309 0.540 124.17 0.25
Table 3: Comparison to optimization-based methods. Our method performs competitively with state-of-the-art optimization-based methods, while reducing the inference cost $300\times$.
[…]